OpenStudy (anonymous):

Integrate sin^2(x)cos^4(x) dx

OpenStudy (anonymous):

use 1-cos 2x=2 sin^2 x and 1+cos2x =2cos ^x to convert them into linear terms

OpenStudy (anonymous):

\[\int\limits_{}^{}(sinxcosx)^2\cos^2dx=\int\limits_{}^{}(\frac{ 1 }{ 2 }\sin2x)^2(\frac{ 1+\cos2x }{ 2 }) dx\]

OpenStudy (anonymous):

i used double angle formula , but im stuck

OpenStudy (anonymous):

and what about the one that i mentioned above??

OpenStudy (anonymous):

yeah, thts double angle right? its the same bt idk how to proceed with this

OpenStudy (anonymous):

when u'll use those one only cos terms will be there.....and they are easy to deal in comparison to sin x.cos x terms